OpenStudy (anonymous):

16x - 8x - 12 = 24 + 20 A. x = 7 B. x = 5 C. x = 4 D. x = 6

OpenStudy (anonymous):

what do you think

OpenStudy (danjs):

a*x + b*x = (a+b)*x

OpenStudy (danjs):

add the coefficients for each like powers of the variable, and solve

OpenStudy (danjs):

(16 - 8)*x = 24 + 20 + 12

OpenStudy (anonymous):

ok so